OpenStudy (sierra_grey):

how do I do this; Express the ratio as a fraction in simplest form. 18 : 12

OpenStudy (shamil98):

\[\huge \frac{ 18 }{ 12 } -->\frac{ 9 }{ 6 } --> \frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

same as reducing \[\frac{18}{12}\]divide top and bottom by the common factor of 6

OpenStudy (shamil98):

Take out the common factors like so. 18/12 both can be reduced by 2, you get 9/6, they can be reduced 3 , you get 3/2

OpenStudy (calculusxy):

An easy way can be to find the GCF of 18 and 12. Factors of 18- 1,2,3,6,9,18 Factors of 12-1,2,3,4,6,12 GCF is 6. You can now divide 18/6=3 and 12/6=2. Now you get the fraction of 3/2.

OpenStudy (sierra_grey):

thanks to everyone! im not sure on who should get the best response ._. its very clear on how to do it now :)

OpenStudy (anonymous):

easy way is to have a good grasp of the multiplication table